# MT-HCCAR: Multi-Task Deep Learning with Hierarchical Classification and Attention-based Regression for Cloud Property Retrieval

### v1

@@ III-B Problem definition for inverse cloud property retrieval
Our primary objective is to accurately model the base-10 logarithm of COT (throughout the remaining text, COT means this logarithm unless explicitly described as the “original” COT) for pixels labeled as cloudy. We work in log space for COT as this has a more linear relationship with the brightness seen by satellite than the original COT. Simultaneously, the model should be able to accurately classify the cloud phase [cloudy, cloud-free, cloudy-liquid, cloudy-ice] for each pixel in the dataset to aid COT prediction. Therefore, our problem consists of two tasks: 1) a classification task to predict cloud mask and phase, and 2) a regression task to predict COT values. Assuming that the labels are $\bar{C}=$ ‘cloud-free’, $C=$ ‘cloudy’, $CL=$ ‘cloudy-liquid’, and $CI=$‘cloudy-ice’, the details of our study are outlined below.
  • Model input. Our model takes 2-dimensional (2D) data as input that contains a set of sample pixels $S={s_{1},s_{2},\ldots,s_{N}}$, where $N$ is the number of labeled pixels. Each pixel has the following features and target variables.
@@
    • Explanatory features: The input features are represented as $\mathbf{X}=[\mathbf{x_{1}},\mathbf{x_{2}},\ldots,\mathbf{x_{M}}]$, where $M$ is the dimension of the available features. The exact input feature variables include: 1) surface pressure in millibar (mbar), 2) total column water vapor content in millimeters (mm), 3) total column ozone content in the Dobson unit (DU), 4) types of Earth surface with 4 categories (land, snow, desert, and ocean or water), 5) top of atmosphere reflectance at different wavelengths collected by each spaceborne sensor, 6) viewing zenith angle, solar zenith angle, and the relative azimuth angle, in degree.
    • Cloud mask/phase label (denoted as $\mathbf{l^{cls}}$): Cloud mask/phase labels are used to train the classification task of the model. The set of possible label values is $l=\{C,\bar{C},CL,CI\}$.
    • True COT value (denoted as $\mathbf{y^{cot}}$): the values are used to train the regression task of the model.
  • Model output
@@
    • Predicted cloud mask/phase class $\mathbf{\hat{l}^{cls}}$ with probabilities $\mathbf{u}$ of each pixel belonging to each of the four classes.
    • Predicted COT value $\mathbf{\hat{y}^{cot}}$ for each pixel.
    • Model parameters $\mathbf{\beta}$.
Specifically, there are three pieces of background knowledge related to the targeted variables $\mathbf{y^{cot}}$ and $\mathbf{l^{cls}}$ in the dataset for model training.
Knowledge 1: Original COT values are not available for pixels with no cloud cover, so not every pixel has COT value. That is, if ${l^{cls}_{i}=0}$, then $y^{cot}_{i}=N/A$.
Knowledge 2: There is a hierarchical relationship between the labels, as pixels covered by liquid cloud or ice cloud are both cloudy pixels. Thus, $CL$ and $CI$ can be two subclasses of label $C=$‘cloudy’.
Knowledge 3: Each pixel is either a cloudy pixel or a cloud-free pixel, and the cloudy pixel is either classified as a cloudy-liquid pixel or a cloudy-ice pixel. There is no coexistence between liquid cloud and ice cloud in our data. That is, each pixel only be labeled as $\bar{C}$, or as $C$ together with one of $\{CL,CI\}$.
With the input features $\mathbf{X}$ and true values of the target variables $[\mathbf{y^{cot}},\mathbf{l^{cls}}]$, we aim to learn a model $\mathcal{M}(\beta)$ such that $[\mathbf{\hat{y}^{cot}},\mathbf{\hat{l}}^{cls}]=\mathcal{M}(\mathbf{X|\beta})$.
IV MT-HCCAR Model
@@
Figure 2: MT-HCCAR: an end-to-end multi-task learning model with hierarchical classification (HC) and cross attention assisted regression (CAR). The HC sub-network consists of the cloud masking module $C_{Mask}(\cdot)$ and the cloud phase classification module $C_{Phase}(\cdot)$. The CAR sub-network consists of the auxiliary coarse classification module $C_{Aux}(\cdot)$, the cross attention module $A(\cdot)$, and the regression module $R(\cdot)$. On the right is the structure of $A(\cdot)$.
Our research objective involves the training of deep learning models to accomplish two tasks: 1) the classification of cloud mask and phase for each pixel based on its reflectance values, and 2) the subsequent prediction of Cloud Optical Thickness (COT) values for pixels classified as cloudy. To address these objectives, we propose an end-to-end Multi-Task Learning (MTL) model, denoted as MT-HCCAR, which integrates a Hierarchical Classification network (HC) and a Classification-assisted with Attention-based Regression network (CAR). Illustrated in Figure 2, the model commences with an encoder, and three branches including Decoder, HC, and CAR follow. Subsequent sections will expound on the individual modules within MT-HCCAR.
IV-A Encoder-decoder sub-network
@@
The shared encoder is formed by the first three dense layers, each of which is followed by a ReLU activation function. This feature extractor is a wide-to-narrow structure where the number of kernels at each layer decreases gradually. The decoder and reconstruction branch form a regularized autoencoder which encourages the model to learn features in the aid of the classification task and regression task.
IV-B Hierarchical classification (HC) sub-network
Utilizing the shared parameters acquired by the encoder-decoder sub-network, MT-HCCAR incorporates a hierarchical classification sub-network. This sub-network comprises a cloud mask classification module and a cloud phase classification module, facilitating cloud mask and phase predictions. To enhance the physical interpretability of the model, the architecture is crafted to mirror human cognitive processes in understanding cloud labels. Fundamentally, to predict cloud phase and Cloud Optical Thickness (COT) values from an Earth science perspective, the model must discern between cloudy and non-cloudy pixels before further classifying liquid-phase and ice-phase pixels and predicting COT values for cloudy pixels.
Hierarchical classification has the capability to categorize instances according to label levels, forming a tree-like structure where each label functions as a node on the tree[39]. The HC network used in MT-HCCAR consists of two classifiers $C_{Mask}(\cdot)$ and ${C_{Phase}}(\cdot)$, the first of which distinguishes cloudy pixels from the cloud-free pixels, and the output is uncertainties for the two labels $[\mathbf{\hat{u}^{C}},\mathbf{\hat{u}^{\bar{C}}}]$. The output uncertainties are binarized to $[\mathbf{\hat{l}^{C}},\mathbf{\hat{l}^{\bar{C}}}]$ by a threshold of 0.5. Before the beginning of the second classifier ${C_{Phase}}(\cdot)$ input feature $\mathbf{X}$ is multiplied by the predicted cloud label, $\mathbf{\hat{y}^{C}}$, to filter out cloud-free pixels. Then ${C_{Phase}}(\cdot)$ classifies liquid-cloud and ice-cloud and the sigmoid uncertainties $[\mathbf{\hat{u}^{CL}},\mathbf{\hat{u}^{CI}}]$ are transformed in the same way to binary cloud-phase labels $[\mathbf{\hat{l}^{CL}},\mathbf{\hat{l}^{CI}}]$.
IV-C Classification assisted regression sub-network based on cross attention mechanism (CAR)
Taking inputs from the encoded feature set and cloud mask classification results as input, our regression sub-network is used for further downstream regression prediction of a cloudy pixel’s COT value. Two novel efforts were taken to improve COT prediction accuracy. First, instead of a direct regression module, an auxiliary coarse classification module is added to predict which sub-range of COT each pixel falls into. Second, a residual-based cross-attention mechanism inspired by [40] is developed to enable the regression module and the auxiliary classification module to share relevant correlations and insights.
@@ IV-C Classification assisted regression sub-network based on cross attention mechanism (CAR)
Our auxiliary coarse classification involves discretizing continuous values into coarse groups, serving as a preprocessing step before regression to align pixels with similar characteristics in the feature space. In MT-HCCAR, an auxiliary classifier is employed to categorize continuous COT values into three distinct levels: thin cloud, moderate cloud, and thick cloud. Specifically, we define thin cloud pixels as those with logarithmic COT values within the range of [-1.5, 0], moderate cloud in [0, 1], and thick cloud in [1, 2.5]. Given the analogous nature of the tasks in COT coarse classification and COT regression, the features extracted by the auxiliary classifier are integrated with those obtained from the regression network.
@@
To learn the joint features more effectively, we utilize cross-attention layers to enhance the integration of features from both regression layers and coarse classification layers, facilitating deeper feature learning and more cohesive feature amalgamation. Based on the cross-attention mechanism, adding a residual connection guarantees that the module has stable output and related research such as Wang et al. [41] propose a non-local operation network using a residual connection to insert blocks to the network. By using the residual connection in our module $A(\cdot)$, the output is
$A(\vartheta_{1},\vartheta_{2})=W_{z}y_{A}(\vartheta_{1},\vartheta_{2})+\vartheta_{1}$ (2)
, where as shown in Figure 2 $\vartheta_{1}$ is the feature from the regression network, $\vartheta_{2}$ is the feature from the auxiliary classification module, $W_{z}$ is the weight matrix to map $y_{A}$ to $\vartheta_{2}$, and $y_{A}$ denotes the computation within cross-attention mechanism. The calculation of $y_{A}$ involves three variables: a query ($Q$), a key ($K$) and a value ($V$)[40]. They are calculated based on corresponding weights matrices $W_{Q}$, $W_{K}$, $W_{V}$ learned through model training. $Q=W_{Q}\vartheta_{2}$, $K=W_{K}\vartheta_{1}$, and $V=W_{V}\vartheta_{1}$. Then the $y_{A}$ in Equation 2 is:
$y_{A}=Softmax(QK^{T})V$ (3)
IV-D Model training of MT-HCCAR
@@
The loss function $\mathcal{L}_{MT-HCCAR}$ for training MT-HCCAR model is formulated as the weighted sum of four components: a hierarchical classification loss $\mathcal{L}_{HC}$, a regression loss $\mathcal{L}_{CAR}$, a reconstruction loss $\mathcal{L}_{Rec}$, and a Lasso regularization loss $\mathcal{L}_{Lasso}$. That is
$\mathcal{L}=\mathcal{L}_{HC}+\mathcal{L}_{CAR}+\mathcal{L}_{Rec}+\mathcal{L}_{Lasso}$ (4)
The four loss components are calculated using different rules. The reconstruction loss $\mathcal{L}_{Rec}$ describes the difference between the input features $\mathbf{X}$ and the reconstruction of the input generated through the encoder and the decoder. It is Mean Square Error (MSE) between $\mathbf{D}(\mathbf{X})=\mathbf{\hat{X}}$ and $\mathbf{X}$:
$\mathcal{L}_{Rec}=\sum_{i=1}^{n}(\mathbf{X_{i}}-\mathbf{\hat{X_{i}}})^{2}$ (5)
The Lasso regularization loss component $loss_{l1}$ is an additional penalty to regularize the training process by increasing the sparsity of the model with $l1$ regularization loss:
$\mathcal{L}_{Lasso}=\lambda\sum_{p=1}^{P}|\beta_{p}|$ (6)
The classification loss and regression loss are a weighted sum of losses from the supplement modules. The HC loss is the summation of Binary Cross Entropy loss from cloud masking module $C_{Mask}(\cdot)$ and cloud phase classification module $C_{Phase}(\cdot)$.
$\mathcal{L}_{HC}=\mathcal{L}_{C_{Mask}}+\mathcal{L}_{C_{Phase}}$ (7)
@@
$\mathcal{L}_{C_{Mask}}=-\frac{1}{N}\sum_{i=1}^{N}l^{C}_{i}\cdot\log(\hat{u}^{C}_{i})+l^{\bar{C}}_{i}\cdot\log(\hat{u}^{\bar{C}}_{i})$ (8)
$\mathcal{L}_{C_{Phase}}=-\frac{1}{N}\sum_{i=1}^{N}\hat{u}^{C}_{i}l^{CL}_{i}\cdot\log(\hat{u}^{C}_{i}\hat{u}^{CL}_{i})+\hat{u}^{C}_{i}l^{CI}_{i}\cdot\log(\hat{u}^{C}_{i}\hat{u}^{CI}_{i})$ (9)
The CAR loss $\mathcal{L}_{CAR}$ is the summation of Cross Entropy loss from the auxiliary classifier $C_{Aux}(\cdot)$ and $l1$ loss from regression.
$\mathcal{L}_{CAR}=\mathcal{L}_{R}+\mathcal{L}_{C_{Aux}}$
@@
We have predicted COT values and $\hat{y}^{cot}$ and sigmoid output of thickness group uncertainties $[\hat{u}^{thin},\hat{u}^{mod},\hat{u}^{thick}]$ from auxiliary classifier, then the regression loss can be formulated as:
$\mathcal{L}_{R}=\sum_{i=1}^{N}|{y}^{cot}_{i}-\hat{y}^{cot}_{i}|$ (10)
$\mathcal{L}_{C_{Aux}}=-\sum_{i=1}^{N}\sum_{c\in[thin,mod,thick]}l^{c}_{i}\ln{\hat{u}^{c}_{i}}$ (11)
V Experiments
This section presents the comparison with baseline methods and ablation studies, where models are trained by the OCI simulated dataset. Section V-A presents the experimental setup for evaluations. Subsequently, Section V-B introduces the evaluation metrics. Following this, Section V-C and Section V-D undertake a comparative analysis between our model and two baseline methods, evaluating performance in both classification and regression tasks. Ablation experiments are then conducted to discern the significance of the modules employed in our model from an earth science perspective.
V-A Experiment setting
The dataset comprises satellite data simulation of instruments OCI, VIIRS and ABI. The three datasets encompassing $N=250,000$ instances. We split data into 62.5%, 22.5%, and 10.0% for training, validation, and test sets, respectively.
To evaluate our model’s performance, we compare MT-HCCAR with six other models on the same dataset OCI. The first two are baseline methods derived from prior studies. The baseline methods are selected from related work with similar task objectives and data formats. Because our dataset contains independent pixels, the baseline methods that can be adapted to our experiments are:
  • Chen et al. (2018): A Multi-layer Perceptron (MLP) based method proposed by Chen et al.[42] for cloud property retrieval. We use 1 hidden layer with 10 nodes, which are the same as the authors used.
  • Liu et al. (2022): A random forest (RF) based method implemented by Liu et al.[8] for cloud detection. We apply one RF to cloud masking and the same RF to cloud phase classification, respectively. The parameters involved are $ntrees=100$ and $mdepth=10$, which are the best parameters for RF found by the authors.

### v2

@@ 3.2 Cloud Property Retrieval
Our primary objective is to accurately model the base-10 logarithm of COT (throughout the remaining text, COT means this logarithm unless explicitly described as the “original” COT) for pixels labeled as cloudy. We work in log space for COT as this has a more linear relationship with the brightness seen by satellite than the original COT. Simultaneously, the model should be able to accurately classify the cloud phase [cloudy, cloud-free, cloudy-liquid, cloudy-ice] for each pixel in the dataset to aid COT prediction. Therefore, our problem consists of two tasks: 1) a classification task to predict cloud mask and phase, and 2) a regression task to predict COT values. Assuming that the labels are $\bar{C}=$ ‘cloud-free’, $C=$ ‘cloudy’, $CL=$ ‘cloudy-liquid’, $CI=$‘cloudy-ice’, and bold notations indicate arrays, the details of our study are outlined below.
@@
Model input. 1) Input features are represented as $\mathbf{X}=[\mathbf{x_{1}},\mathbf{x_{2}},\ldots,\mathbf{x_{M}}]$, where $M$ is the dimension of the available features. The exact input feature variables include: i) surface pressure in millibar (mbar), ii) total column water vapor content in millimeters (mm), iii) total column ozone content in Dobson units (DU), iv) types of Earth surface with 4 categories (land, snow, desert, and ocean/water), v) top of atmosphere reflectance at different wavelengths collected by each spaceborne sensor, vi) viewing zenith angle, solar zenith angle, and the relative azimuth angle, in degree. 2) Cloud mask/phase labels $\mathbf{l^{cls}}$ are used to train the classification task of the model. The set of possible label values is $l=\{C,\bar{C},CL,CI\}$. There is a hierarchical relationship between the labels, as pixels covered by liquid cloud or ice cloud are both cloudy pixels. Also, there is no coexistence between liquid cloud and ice cloud in our data. Thus, $CL$ and $CI$ can be two subclasses of label $C=$‘cloudy’. 3) True COT value $\mathbf{y^{cot}}$ are used to train the regression task of the model. COT values are not available for pixels with no cloud cover. That is, if ${l^{cls}_{i}=0}$, then $y^{cot}_{i}=N/A$.
@@
Model output. 1) Predicted cloud mask/phase class $\mathbf{\hat{l}^{cls}}$ with probabilities $\mathbf{u}$ of each pixel belonging to each of the four classes. 2) Predicted COT value $\mathbf{\hat{y}^{cot}}$ for each pixel. 3) Model architecture $\mathcal{M}$ with parameters $\mathbf{\beta}$. The predicted values are generated by the model, which is $[\mathbf{\hat{y}^{cot}},\mathbf{\hat{l}}^{cls}]=\mathcal{M}(\mathbf{X|\beta})$.
4 MT-HCCAR Model
@@
Figure 2: MT-HCCAR: an end-to-end MTL model with hierarchical classification (HC) and cross attention assisted regression (CAR). The HC sub-network comprises the cloud masking module $C_{Mask}$ and the cloud phase classification module $C_{Phase}$. The CAR sub-network consists of the auxiliary coarse classification module $C_{Aux}$, the cross attention module $A$, and the regression module $R$. On the right is the structure of $A$.
Our objective is to train a deep learning model to accomplish two tasks: 1) the classification of cloud mask and phase for each pixel based on its reflectance values, and 2) the subsequent prediction of COT values for pixels classified as cloudy. Toward this objective, our proposed MT-HCCAR model is illustrated in Figure 2 and the subsections below will each of the modules.
4.1 Encoder-Decoder Sub-Network
@@
The shared encoder is formed by the first three dense layers, each of which is followed by a ReLU activation function. This feature extractor is a wide-to-narrow structure where the number of kernels at each layer decreases gradually. The decoder and reconstruction branch is narrow-to-wide and further improves the performance of the encoder to learn features for different tasks.
4.2 Hierarchical Classification (HC) Sub-Network
Utilizing the shared parameters acquired by the encoder-decoder sub-network, MT-HCCAR incorporates a hierarchical classification sub-network. This sub-network comprises a cloud mask classification module and a cloud phase classification module, facilitating cloud mask and phase predictions. To enhance the physical interpretability of the model, the architecture is crafted to mirror human cognitive processes in understanding cloud labels. Fundamentally, to predict cloud phase and COT values from an Earth science perspective, the model must discern between cloudy and non-cloudy pixels before further classifying liquid-phase and ice-phase pixels and predicting COT values for cloudy pixels.
Hierarchical classification has the capability to categorize instances according to label levels, forming a tree-like structure where each label functions as a node on the tree [18]. The HC network used in MT-HCCAR consists of two classifiers $C_{Mask}$ and ${C_{Phase}}$, the first of which distinguishes cloudy pixels from the cloud-free pixels, and the output is uncertainties for the two labels $[\mathbf{\hat{u}^{C}},\mathbf{\hat{u}^{\bar{C}}}]$. The output uncertainties are binarized to $[\mathbf{\hat{l}^{C}},\mathbf{\hat{l}^{\bar{C}}}]$ by a threshold of 0.5. Before the beginning of the second classifier ${C_{Phase}}$ input feature $\mathbf{X}$ is multiplied by the predicted cloud label, $\mathbf{\hat{y}^{C}}$, to filter out cloud-free pixels. Then ${C_{Phase}}$ classifies liquid-cloud and ice-cloud and the sigmoid uncertainties $[\mathbf{\hat{u}^{CL}},\mathbf{\hat{u}^{CI}}]$ are transformed in the same way to binary cloud-phase labels $[\mathbf{\hat{l}^{CL}},\mathbf{\hat{l}^{CI}}]$.
4.3 Classification Assisted Regression Sub-Network based on Cross Attention Mechanism (CAR)
Taking inputs from the encoded feature set and cloud mask classification results as input, our regression sub-network CAR is used for further downstream regression prediction of a cloudy pixel’s COT value. Two novel efforts were taken to improve COT prediction accuracy. First, instead of a direct regression module, an auxiliary coarse classification module is added to predict which sub-range of COT each pixel falls into. Second, a residual-based cross-attention mechanism inspired by [23] is developed to enable the regression module and the auxiliary classification module to share relevant correlations and insights.
@@ 4.3 Classification Assisted Regression Sub-Network based on Cross Attention Mechanism (CAR)
Our auxiliary coarse classification involves discretizing continuous values into coarse groups, serving as a preprocessing step before regression to align pixels with similar characteristics in the feature space. Moreover, the auxiliary classifier assigns greater importance to the regression task during model training to get accurate COT predictions. In MT-HCCAR, an auxiliary classifier is employed to categorize continuous COT values into three distinct levels: thin cloud, moderate cloud, and thick cloud. Specifically, we define thin cloud pixels as those with logarithmic COT values within the range of [-1.5, 0], moderate cloud in [0, 1], and thick cloud in [1.0, 2.5]. Given the analogous nature of the tasks in COT coarse classification and COT regression, the features extracted by the auxiliary classifier are integrated with those obtained from the regression network.
@@
To learn the joint features more effectively, we utilize cross-attention layers to enhance the integration of features from both regression layers and coarse classification layers, facilitating deeper feature learning and more cohesive feature amalgamation. Based on the cross-attention mechanism, adding a residual connection guarantees that the module has a stable output and related research such as Wang et al. [27] propose a non-local operation network using a residual connection to insert blocks to the network. By using the residual connection in our module $A$, the output is $A(\vartheta_{1},\vartheta_{2})=W_{z}y_{A}(\vartheta_{1},\vartheta_{2})+\vartheta_{1}$, where as shown in Figure 2 $\vartheta_{1}$ is the feature from the regression network, $\vartheta_{2}$ is the feature from the auxiliary classification module, $W_{z}$ is the weight matrix to map $y_{A}$ to $\vartheta_{2}$, and $y_{A}$ denotes the computation within cross-attention mechanism. The calculation of $y_{A}$ involves three variables: a query ($Q$), a key ($K$) and a value ($V$[23]. They are calculated based on corresponding weights matrices $W_{Q}$, $W_{K}$, $W_{V}$ learned through model training. $Q=W_{Q}\vartheta_{2}$, $K=W_{K}\vartheta_{1}$, and $V=W_{V}\vartheta_{1}$. Then the $y_{A}$ is $y_{A}=Softmax(QK^{T})V$.
4.4 Model Training of MT-HCCAR
@@
The loss function $\mathcal{L}_{MT-HCCAR}$ for training MT-HCCAR model is formulated as the weighted sum of four components: a hierarchical classification loss $\mathcal{L}_{HC}$, a regression loss $\mathcal{L}_{CAR}$, a reconstruction loss $\mathcal{L}_{Rec}$, and a Lasso regularization loss $\mathcal{L}_{Lasso}$. That is: $\mathcal{L}=\mathcal{L}_{HC}+\mathcal{L}_{CAR}+\mathcal{L}_{Rec}+\mathcal{L}_{Lasso}$. The four loss components in the loss function are calculated using different rules.
@@
The other two components, HC loss and CAR loss, are both weighted sums of losses from the supplement modules. The HC loss is the summation of Binary Cross Entropy loss from cloud masking module $C_{Mask}$ and cloud phase classification module $C_{Phase}$. The HC loss function is $\mathcal{L}_{HC}=\mathcal{L}_{C_{Mask}}+\mathcal{L}_{C_{Phase}}$. We have a sigmoid output of cloudy and non-cloud uncertainties $[\hat{u}^{C},\hat{u}^{\bar{C}}]$ from cloud mask classifier $C_{Mask}$ and sigmoid output of liquid cloud and ice cloud uncertainties $[\hat{u}^{CL},\hat{u}^{CI}]$ from cloud phase classifier $C_{Phase}$, then for the HC loss, $\mathcal{L}_{C_{Mask}}$ and $\mathcal{L}_{C_{Phase}}$ are calculated as $\mathcal{L}_{C_{Mask}}=-\frac{1}{N}\sum_{i=1}^{N}l^{C}_{i}\cdot\log(\hat{u}^{C}_{i})+l^{\bar{C}}_{i}\cdot\log(\hat{u}^{\bar{C}}_{i})$ and $\mathcal{L}_{C_{Phase}}=-\frac{1}{N}\sum_{i=1}^{N}\hat{u}^{C}_{i}l^{CL}_{i}\cdot\log(\hat{u}^{C}_{i}\hat{u}^{CL}_{i})+\hat{u}^{C}_{i}l^{CI}_{i}\cdot\log(\hat{u}^{C}_{i}\hat{u}^{CI}_{i})$.
@@
The CAR loss $\mathcal{L}_{CAR}$ is the summation of Cross Entropy loss from the auxiliary classifier $C_{Aux}$ and $l1$ loss from regression. That is, $\mathcal{L}_{CAR}=\mathcal{L}_{R}+\mathcal{L}_{C_{Aux}}$. We have predicted COT values and $\hat{y}^{cot}$ and sigmoid output of thickness group uncertainties $[\hat{u}^{thin},\hat{u}^{mod},\hat{u}^{thick}]$ from the auxiliary classifier, then the two components are $\mathcal{L}_{R}=\sum_{i=1}^{N}|{y}^{cot}_{i}-\hat{y}^{cot}_{i}|$ and $\mathcal{L}_{C_{Aux}}=-\sum_{i=1}^{N}\sum_{c\in[thin,mod,thick]}l^{c}_{i}\ln{\hat{u}^{c}_{i}}$.
5 Experiments
To evaluate our model’s performance, we compare MT-HCCAR with baseline methods and conduct an ablation study across OCI, VIIRS, and ABI datasets.
Given that our dataset comprises independent pixels, we choose two baseline methods from prior research with comparable task objectives and data formats for our experiments: 1) Chen et al. (2018) [8]: An MLP based method for cloud property retrieval. We use 1 hidden layer with 10 nodes, which are the same as the authors used. 2) Liu et al. (2022) [17]: An RF based method for cloud detection. We apply one RF to cloud masking and the same RF to cloud phase classification, respectively. The parameters involved in the method are $ntrees=100$ and $mdepth=10$.